(a) A nurse has two solutions that contain different concentrations of a certain medication. One is a 12.5% concentration and the other is a 5% concentration. How many cubic centimeters of each should she mix to obtain 20 cubic centimeters of an 8% concentration?

Answer by stanbon(75887)

(a) A nurse has two solutions that contain different concentrations of a certain medication. One is a 12.5% concentration and the other is a 5% concentration. How many cubic centimeters of each should she mix to obtain 20 cubic centimeters of an 8% concentration?

-------------
Equation::
active + active = active
------
0.125x + 0.05(20-x) = 0.08*20
------
125x + 50(20-x) = 80*20
125x + 50*20 - 50x = 80*20
75x = 30*20
x = 8 cc (amt. of 12.5% solution needed)
20-x = 12 cc (amt. of 8% solution needed)
----------------
Cheers,
Stan H.
==================